Fermi National Laboratory


Accelerator Update for December 14, 2005 - December 16, 2005

Wednesday December 14

The day shift began with Operations monitoring store 4549 and a Recycler stash, with the Antiproton Source (Pbar) stacking, with NuMI off due to a Horn chiller problem, and with MiniBooNE and the Meson MCenter experiment taking beam.

At 8:26 AM, a Booster expert asked to have a RF station (BRF19) turned off.

Operations established beam to the Meson MTest experiment at 9:39 AM. They reported that they were done with beam for the day at 12:36 PM.

Operations began preparing Recycler to stash antiprotons at 2:50 PM. The stash ended at 3:59 PM, but when Pbar tried to resume stacking, the Main Injector (MI) began having a problem.

At 4:15 PM, the Tevatron (TeV) suffered a sector E1 quench due to a B11 separator spark. Operators and the cryo system technicians began the TeV cryo system recovery.

Operations prepared the Collider Detector at Fermilab (CDF) and D-Zero (D0) collision halls for access at 4:41 PM. CDF went into access at 5:36 PM.

At 5:41 PM, Operations reported a problem with an A1 transfer line power supply (I:V901). An operator called in power supply experts.

Operations began the TeV turn on at 6:50 PM.

MTest decided that they weren’t actually through with beam for the day and resumed taking beam at 7:23 PM.

Operations put the TeV through a dry squeeze at 8:59 PM.

CDF came out of access at 8:30 PM.

MTest again reported at 9:24 PM that they really had enough beam for the day.

Power supply techs reported at 10:04 PM that they had finished their repair of I:V901.

The Recycler began preparing to stash at 11:37 PM.

Thursday December 15

The midnight shift began with the Recycler stashing, with NuMI off due to a horn chiller problem, and with MiniBooNE taking beam.

The stash ended and Pbar resumed stacking at 12:15 AM.

Operations reported at 1:06 AM that MiniBooNE had been tripping off regularly due to beam position problems. Operators investigated this problem throughout most of the midnight shift.

At 3:02 AM, Operations reported that Recycler RF problems were holding up the TeV from going into shot setup. An expert came in to investigate.

At 3:46 AM, Operations reported a Pbar LCW leak. Operators had been adding about 20 gallons of water every 30 minutes. Operators searched for signs of a leak.

At 5:29 AM, Operations reported that during the process of Recycler trying to fix their RF problems, they lost all of their stash due to a damper problem. Experts came in and worked on the problems.

The Recycler began preparing to stash at 7:58 AM.

At 8:10 AM, Operations reported that the B1 separator sparked.

The stash ended at 8:55 AM. MCenter resumed taking beam at 9:12 AM.

Operations began preparing Pbar for access at 9:17 AM.

Operators found a LCW leak in the AP1 transfer line at 9:56 AM.

Water technicians had the AP1 LCW leak repaired by 11 AM. Pbar resumed stacking.

At 11:19 AM, Operations reported that TeV experts decided to condition the separators for twenty-four hours more.

At 1:22 PM, Pbar reported having controls problems. Experts investigated.

D0 accessed their collision hall at 1:22 PM and came out about 1.5 hours later.

At 4:57 PM, Electron Cooling experts reported having trouble with the electron beam. They had the problem resolved an hour later.

The Recycler began stashing at 7:23 PM. The stash ended at 8:01 PM.

Pbar and MCenter resumed taking beam by 8:17 PM.

Friday December 16

The midnight shift began with the TeV separators in a conditioning mode, with Operations monitoring the stash, with Pbar stacking, with NuMI off due to Horn chiller problems, and with MiniBooNE and MCenter taking beam.

The Recycler began preparing to stash at 5:50 AM. The stash ended. Pbar and MCenter resumed taking beam at 6:59 AM.

The Plans for Friday and the Weekend

The plans for today are to establish a store later this morning. After the store goes in, operators will access MI and look for the NuMI LCW leak. The plans for the weekend are to stack, stash, and store. There is a hope that we will have NuMI back on line this weekend.

Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now.

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.



last modified 12/16/2005   email Fermilab